OpenStudy (anonymous):

Which expressions, when evaluated with y = –7, are positive in value? Choose all answers that are correct. A. 18 + 2y B. –3y + (–17) C. –2y – 23 D. –30 – 4y

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

You can solve all of them by subbing the value of y with -7 OR an easier way to do it is count the signs two negatives make 1 positive so think of it as this: Odd numbers of negative signs=negative solution Even number of negative signs=postitve solution Let's start with A \[\huge~\rm~\bf~18 + 2(\color{red}{-7})\]

OpenStudy (anonymous):

11

pooja195 (pooja195):

-7 * 2=?

OpenStudy (anonymous):

-14

pooja195 (pooja195):

18-14=?

OpenStudy (anonymous):

4

pooja195 (pooja195):

Good so A would be one of our answers... Next lets try B –3(-7) + (–17) -3 * -7=?

OpenStudy (anonymous):

10 and 21

pooja195 (pooja195):

We are multiplying so just 21 21-17=?

OpenStudy (anonymous):

4

pooja195 (pooja195):

Good so we have A and B as our answers Lets try C -2(-7)-23=?

OpenStudy (anonymous):

-23

pooja195 (pooja195):

-14-23 will result in a negative solution so C is incorrect lets try D –30 – 4(-7) -4*-7=?

OpenStudy (anonymous):

28

pooja195 (pooja195):

good -30+28=?

OpenStudy (anonymous):

-2

pooja195 (pooja195):

Good so C and D can't be answers. Therefore our final answers are...?

OpenStudy (anonymous):

a and b thx :)

pooja195 (pooja195):

yw ^_^